OpenStudy (mattyboyy):

Which ordered pair is a solution of the equation 4x − 2y = 16? A) (-5,-3) B) (-5,3) C) (6,-4) D) (6,4)

OpenStudy (anonymous):

put in the values into the equation 4(6)-2(4)=24-8=16

OpenStudy (mattyboyy):

That equal 16?

OpenStudy (anonymous):

yes

OpenStudy (mattyboyy):

So it would be D then?

OpenStudy (anonymous):

yes

OpenStudy (mattyboyy):

Thanks